OpenStudy (anonymous):

Isabelle made a cylindrical hat for a costume. If the hat is 339 cubic inches and a diameter of 6 inches, what is the approximate height of the hat? Use 3 for pi.

OpenStudy (anonymous):

@timo86m

OpenStudy (anonymous):

please.

OpenStudy (anonymous):

339=pi r^2*h 339=3*3^2*h solve for h

OpenStudy (anonymous):

hmm.alright.

OpenStudy (anonymous):

Approximately 4.18?

OpenStudy (anonymous):

Yes:) thank you.

OpenStudy (anonymous):

h = 12.555555

OpenStudy (anonymous):

Is the website right?or you?

OpenStudy (anonymous):

I will go ahead and post my next q.

OpenStudy (anonymous):

339=3*3^2*h is h=339/(3*3^2)

OpenStudy (anonymous):

1017 in^3 ?

OpenStudy (anonymous):

ooooh. u were right.

OpenStudy (anonymous):

it IS 12.5

OpenStudy (anonymous):

i did 339/27. it was right. thanks,

OpenStudy (anonymous):

yup :)